Propane (C3H8) is a fuel commonly burned for needs such as home heating and cooking food. When propane burns, it reacts with oxygen (O2). The products of the reaction are carbon dioxide (CO2) and water (H2O). If 44.1 g of propane react completely with 160.0 g of oxygen, 132.0 g of carbon dioxide are formed. What is the mass of the other product, water, formed in this reaction?



1
Expert's answer
2022-01-24T15:00:02-0500

Balanced equation for the reaction;

C3H8 + 5 O2 = 3CO2 + 4H2O

molar mass of water= 18g/mol

if 1mole = 18g

4moles=?

= (4 × 18) ÷ 1 =72g

mass of water formed= 72g